# Learned representation-guided diffusion models for large-image generation

### v1

@@ 12 Memory requirements
Training
Method
Maximum
batch size
Training time
per epoch
Ours
(256 $\times$ 256 patches)
100 45
LDM on
1024 $\times$ 1024 images
4 300
@@ 13 Large image generation details
The LDM is applied to patches in the large image with a stride of 16. Using a larger stride leads to tiling artifacts, whereas a smaller stride increases the computational cost without much difference in the synthesized image quality. For each location $i,j$ at which we want to apply the diffusion model, we interpolate the 4-nearest embeddings to get the conditioning $\lambda_{i,j}$. We found that spherical linear interpolation (slerp), weighted by the distance of $i,j$ to the centers of its four neighbors, worked best for interpolating the high-dimensional, normalized SSL embeddings.
@@
For the text-to-large image experiments, we trained an auxiliary diffusion model to sample a $4\times 4$ grid of embeddings given the text conditioning. We used a small convolutional network with residual layers to implement the diffusion model. The timestep conditioning was concatenated to the input grid. The network directly predicted the final embeddings from the conditioning and current noisy embedding grid, instead of predicting the noise added. For TCGA-BRCA and TCGA-CRC, the text conditioning is a single 512-dim Quilt embedding vector. For NAIP, we used a frozen CLIP [35] text encoder to extract features from the text captions and used them as conditioning. For the diffusion process, we used 1000 steps with a linear schedule, as in [16]. Additionally, when sampling embeddings from text for TCGA-BRCA and TCGA-CRC we used negative prompting [25] to further separate the different types of images during generation.
@@ 14 Text-to-large image generation examples
In Fig. 11 we present generated images from the TCGA-BRCA model and the text prompts used in generating them. We borrow the text prompts from the zero-shot classification experiments of [18]. As discussed for the confusion matrix (Fig. 5 in the main text), the vision-language model’s capabilities limit the quality of our results. The model seems to be able to only differentiate between non-malignant / normal and malignant, which is expected since the zero-shot classification accuracy of Quilt on breast cancer images is around 40%. In contrast, for the CRC data where accuracy is around 90%, our text-to-large image generation performs better. In Fig. 12 we present such synthetic samples from TCGA-CRC.
To train the satellite text-to-large image auxiliary diffusion model we generated a synthetic set of image-caption pairs using BLIP [22]. For training, we created a set of 30k large images ($1024\times 1024$ pixels) with 4 captions for each, whereas for the test set, we used a single caption for evaluation. In Fig. 13 we present images from the training and test sets as well as generated samples along with their text prompts. We see that although the training captions are far from perfect, we are able to generate test set images consistent with the prompts used. Even though our training set is tiny, we see interesting generalization capabilities when using ’unusual’ prompts, such as ”a satellite image of a forest with smoke”, where the model tries to add clouds to mimic the ”smoke” seen from a satellite image. This generalization can be attributed to both the expressivity of the SSL embeddings used in synthesizing the images and the usage of a pre-trained CLIP text encoder to interpret the captions.
@@ 15 Pathologist evaluation
We designed a simple user interface where we presented large TCGA-CRC images generated from text prompts and asked an expert pathologist to evaluate them (Fig. 9). The model generated an image using one of two text prompts: ”benign colonic tissue” or ”colon adenocarcinoma”. We asked a pathologist to evaluate by categorizing the images as benign / adenocarcinoma / undecided as well as assigning a realistic / unrealistic label. For a total of 100 images, the final agreement between text prompts and pathologist labels was 89.9%, with 61% of the images marked as realistic. This clearly illustrates the applicability of our proposed method; an auxiliary diffusion model that generates the SSL conditioning from any related modality can be chained with our patch-based diffusion to synthesize coherent large images.

### v2

@@ 12 Memory requirements
Training Method Maximum batch size Training time per epoch
Ours (256 $\times$ 256 patches) 100 45 hr
LDM on 1024 $\times$ 1024 images 4 300 hr
@@
Conditioning Patch FID
None 25.62
ImageNet ViT-B/16 13.29
CLIP [35] 16.07
HIPT [6] 6.98
Table 7: FIDs when using different representations as conditions.
Stride Time/ Crop CLIP
Image FID FID
4 15m 12.66 7.31
8 4m 14.69 7.37
16* 1m 15.51 7.43
32 20s 15.60 8.09
Table 8: Large image generation parameters ablation. By * we denote the stride used in the main text experiments.
13 Using different SSL encoders
We extend the TCGA-BRCA 20×20$\times$20 × model of Table 1 with additional patch-level FID values, obtained by using different embeddings as conditioning (Table 8). The pathology-specific HIPT performs best, suggesting that the domain expressivity of the embedding used as conditioning affects image generation quality. We conjecture that worse patch quality also hurts large image metrics.
@@ 14 Large image generation details
The LDM is applied to patches in the large image with a stride of 16. Using a larger stride leads to tiling artifacts, whereas a smaller stride increases the computational cost without much difference in the synthesized image quality. In Table 8 we provide an ablation study of the large image generation parameters. We synthesize $1024\times 1024$ px images from TCGA-BRCA with different strides, using 50 steps of diffusion, on an NVIDIA RTX 6000, showing that larger strides require fewer forward passes (less time) but produce worse results.
For each location $i,j$ at which we want to apply the diffusion model, we interpolate the 4-nearest embeddings to get the conditioning $\lambda_{i,j}$. We found that spherical linear interpolation (slerp), weighted by the distance of $i,j$ to the centers of its four neighbors, worked best for interpolating the high-dimensional, normalized SSL embeddings.
@@
For the text-to-large image experiments, we trained an auxiliary diffusion model to sample a $4\times 4$ grid of embeddings given the text conditioning. We used a small convolutional network with residual layers to implement the diffusion model. The timestep conditioning was concatenated to the input grid of embeddings. The network directly predicted the final embeddings from the conditioning and current noisy embedding grid, instead of predicting the noise added. For TCGA-BRCA and TCGA-CRC, the text conditioning is a single 512-dim Quilt embedding vector. For NAIP, we used a frozen CLIP [35] text encoder to extract features from the text captions and used them as conditioning. For the diffusion process, we used 1000 steps with a linear schedule, as in [16]. Additionally, when sampling embeddings from text for TCGA-BRCA and TCGA-CRC we used negative prompting [25] to further separate the different types of images during generation.
@@ 15 Text-to-large image generation examples
In Fig. 10 we present generated images from the TCGA-BRCA model and the text prompts used in generating them. We borrow the text prompts from the zero-shot classification experiments of [18]. As discussed for the confusion matrix (Fig. 4), the vision-language model’s capabilities limit the quality of our results. The model seems to be able to only differentiate between non-malignant / normal and malignant, which is expected since the zero-shot classification accuracy of Quilt on breast cancer images is around 40%. In contrast, for the CRC data where accuracy is around 90%, our text-to-large image generation performs better. In Fig. 11 we present such synthetic samples from TCGA-CRC.
To train the satellite text-to-large image auxiliary diffusion model we generated a synthetic set of image-caption pairs using BLIP [22]. For training, we created a set of 30k large images ($1024\times 1024$ pixels) with 4 captions for each, whereas for the test set, we used a single caption for evaluation. In Fig. 12 we present images from the training and test sets as well as generated samples along with their text prompts. We see that although the training captions are far from perfect, we are able to generate test set images consistent with the prompts used. Even though our training set is tiny, we see interesting generalization capabilities when using ’unusual’ prompts, such as "a satellite image of a forest with smoke", where the model tries to add clouds to mimic the "smoke" seen from a satellite image. This generalization can be attributed to both the expressivity of the SSL embeddings used in synthesizing the images and the usage of a pre-trained CLIP text encoder to interpret the captions.
@@ 16 Pathologist evaluation
We designed a simple user interface where we presented large TCGA-CRC images generated from text prompts and asked an expert pathologist to evaluate them (Fig. 8). The model generated an image using one of two text prompts: "benign colonic tissue" or "colon adenocarcinoma". We asked a pathologist to evaluate by categorizing the images as benign / adenocarcinoma / undecided as well as assigning a realistic / unrealistic label. For a total of 100 images, the final agreement between text prompts and pathologist labels was 89.9%, with 61% of the images marked as realistic. This clearly illustrates the applicability of our proposed method; an auxiliary diffusion model that generates the SSL conditioning from any related modality can be chained with our patch-based diffusion to synthesize coherent large images.